Find the value of the following: 
3845 × 5 × 782 + 769 × 25 × 218

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given
3845 × 5 × 782 + 769 × 25 × 218
= 3845 × 5 × 782 + 769 × 5 × 5 × 218
= 3845 × 5 × 782 + 3845 × 5 × 218
= 3845 × 5 × (782 + 218)
= 19225 × 1000
= 19225000

Thanks